Insert a A.M. between 5 and 20.

Grade:Upto college level

2 Answers

vikas askiitian expert
13 years ago

AM(arithemetic mean) = a+b/2

here in this question a = 5 & b = 20 so

AM = 20+5/2 = 25/2

now series is 5 , 25/2 , 20

common difference = 15/2 & first term is 5

 

approve if u like my ans

Ashwin Sinha
13 years ago

Dear Anitha,

Let a=5   &    b=20.

 

  Now common diff.= (b-a)/2 =(20-5)/2=15/2 =d

        So,   A.M. is a+d =5+15/2=25/2.

        Hope it helps!!

      Plz approve!!